# An Initial Exploration of Default Images in Text-to-Image Generation

### v1

@@ Abstract.
The figure shows different default images, resulting from unrelated input prompts. It is clearly visible that these images have a common theme, motive (i.e., subjects), and style.
@@ 1. Introduction
While there is a growing number of studies investigating TTI and the practice of prompt engineering, little attention has been paid to what happens when the model does not (in anthropomorphic terms) understand its inputs. TTI models are trained to always yield an output, even if the input prompt contains terms and concepts not known to the model. Therefore, when such a prompt is given to the TTI model, it may generate what we call default images – images that are seemingly unrelated to the input prompt and often do not match the user’s expectations. One key characteristic of default images is that the same (or very similar) default image may appear when prompted with different, unrelated prompts (see Figure 1 and Figure 3).
@@
In this paper, we present the first preliminary exploration into default images in Midjourney, a popular image generation system. We describe our approach to systematically create a set of input prompts likely triggering default images, and describe our experiments in generating images from these prompts. We then apply affinity diagramming to identify similarities between the generated images, resulting in a set of ten default images. We describe these images and make nine postulates about their characteristics. We further conducted a user study investigating whether default images affect user satisfaction, and how they align with human perception and expectations. Note that due to the vastness of Midjourney’s latent space, this initial investigation is not exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and the importance of studying default images.
@@
  • A description of our systematic approach for generating default images. Due to the vastness of Midjourney’s latent space, this investigation can never be exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and present the results of several ablation studies.
@@ 3.1. What are Default Images?
The defining characteristic of default images is that they are encountered in response to multiple different and—crucially—unrelated prompts. The examples given in Figure 1 and Figure 3 demonstrate how different unrelated prompts trigger very similar default images in Midjourney – images with a specific motive and style that originate from a region in the TTI model’s latent space that is not strongly associated with known concepts. Another defining characteristic of default images is that, while there are slight variations, the default images are remarkably stable in terms of their motive (i.e., subject) and style (cf. Figure 1). Their recurring motive and style makes default images distinctly recognizable.
@@
3.2. Importance of Studying Default Images in Text-to-Image Generation
@@ 4.1. Input Prompts Likely Triggering Default Images
In our initial exploration of the phenomenon of default images, we explored three categories of prompts that could result in a conceptual mismatch between the resulting image and user expectations. The first category includes prompts that are likely to confuse the Midjourney model. This category includes tokens that were likely not part of the model’s training data, such as email addresses, polysemous words, ambiguous topics, and censored words (e.g., political topics). While these prompts are valuable for testing the limits of Midjourney’s models, the probability of default images appearing is too low in our experience. This causes the prompts of this first category to be unsuitable for our systematic approach for generating default images.
The second category involves prompts that are not recognized by the model and lack practical purpose. These prompts include, but are not limited to: contextually misleading phrases (e.g., Einstein’s relativity theology), random character strings and special characters (e.g. aölkjnewjbtn and %), misheard lyrics, and abstract and metaphorical concepts that may not invoke strong imagery (such as absolute ownership and extreme aptitude). While prompts in this category may lead to default images, we excluded them from our analysis due to the limited relevance in real-world applications. These prompts do not reflect how users would typically interact with the TTI model, although they are relevant to for future investigations.
The third category is the most promising to our endeavor, combining both terms not recognized by the model and practical use cases. These prompts are likely used in the real world, but still challenge the model. This includes:
@@
The six categories of prompts emphasize real-world usefulness and they are our focus area and inspiration for creating default images. The low-resource languages were chosen for their language-native words that are not borrowed from other languages and, thus, do not cognate with widely known foreign languages. We focus on Finnish in this study (see Appendix A4.1), leaving Tagalog (Appendix A4.2) to future work. Using this approach, we created a set of prompts, listed in Appendix A. Each of the prompts was checked on haveibeentrained.com, a website that allows to text against the training data used in many TTI models (including LAION (LAION.ai, nd; Schuhmann et al., 2022)). We then generated images with these prompts on Midjourney, as described in the following section.
@@ 4.2. Image Generation Experiments and Ablation Studies
Our experiments were executed on Midjourney’s ’Create’ page using fixed parameters listed in Table 1. For each of the six different categories of inputs A1–A6, we ran 20 trials (with exception of category glitch tokens, for which we ran 10 trials due to limited availability of glitch tokens), introducing slight variations between each run by corrupting or altering words (see the prompts in Appendix A). This resulted in 440 images. The images were then visually analyzed, as described in the following section.
@@ 4.4. Survey Study
In addition to our visual analysis with affinity diagramming, we utilized crowdsourcing on Prolific111www.prolific.com, an online human subject pool, to inquire about people’s perception of default images. In particular, default images may negatively affect user satisfaction when prompting TTI models. To investigate this, we designed an online study with a questionnaire in which four images were presented, including the prompt that created the image. Unknown to the participant, two of the images (Q1 and Q2) had one term in the prompt replaced with a term that caused default images in our experiments. For instance, in Q1, we replaced the term ‘raven’ in the prompt “photoshoot of a woman wearing a white off-shoulder dress with a black raven on top of her head” with the word ’greagoft’. The participant, however, was shown the original prompt and the image resulting from the modified prompt. This setup simulates the surprising mismatch in expectations that an encounter with a default image may produce in practice. For the third question (Q3), we used an image that matches exactly the descriptive prompt, without replacements. In the fourth question (Q4), we used a prompt which is unknown to the model and, thus, the accompanying image is a default image. The four prompts and images are depicted in Figure 6. Participants were asked to rate their satisfaction with each image, given the prompt, on a scale from 1 (Strongly Dissatisfied) to 7 (Strongly Satisfied). The reward to participants was £6.29 per hour, and the median time to complete the survey was 3 min and 49 seconds.
@@ 5.2.1. Effects of seed values
We visually analyzed how different seed values affect the default images. Using the different seeds with the same prompt resulted in different default images, partially sharing some of the motives (see left images in Figure 5A). While the premise is still the same, one might not associate these images with each other. Thus, the best way to generate default images is by having a fixed seed value for all image generation runs. The images on the right of Figure 5A show how common elements may occur in default images. The commonalities in the default images included portraits of women and animals that were mostly birds.
Figure 5. Ablation studies: (A) Default images generated with different seed values (three images on the left) and image similarities with variations in both seed values and prompts (on the right). (B) The default images on the left demonstrate how known terms (computer, browsing) overpower the unknown terms (in bold), producing images that are not default images. The two images on the right demonstrate how motifs from default images can be created in specific contexts. (C) Default images generated using the prompt “hevonen” and seed 123456 with Midjourney model version 6.1, 6.0, 5.2, and 5.1. Similarities in default images point to similarities in model training data. The four different images generated in response to one prompt (on the left) are the result of Midjourney’s latent space sampling and internal image diversity mechanisms.
@@ 5.3.2. User study results
Figure 6 shows that unknown terms potentially triggering default images may negatively affect user satisfaction, but not in all cases. In Q1, the effect of replacing ‘raven’ with ‘greagoft’ resulted in only subtle changes to the resulting image. The image appeared almost as if a raven could be there. The participants’ satisfaction, therefore, was high ($Mean=5,2$, $SD=1,8$). In Q2, where ‘leaves and flowers’ was replaced with ‘greagoft’, the divergence between prompt and resulting image was more noticeable and satisfaction was lower ($Mean=2,7$, $SD=1,9$). A Fisher’s exact test indicates that The way participants perceived the change (subtle vs. noticeable) is significantly related to their satisfaction level ($OR=0.074$, $p<0.05$). The odds ratio of 0.074 indicates that dissatisfaction (1–3 on the Likert scale) is much less likely when the change is subtle compared to noticeable. In other words, participants were significantly more likely to be satisfied (5–7) when the change in the image was subtle (Q1) rather than noticeable (Q2).
For the third question (Q3), we did not switch or omit any words in the prompt, and the image was the exact result of the prompt. Participants were satisfied with the resulting image ($Mean=4,47$, $SD=1,87$). In the fourth question (Q4), we used a prompt known to trigger default images. Participants showed particularly strong dissatisfaction with this image ($Mean=2.6$, $SD=2.0$). Again, there was a significant difference in user satisfaction between Q3 (representing an exact match between prompt and image) and Q4 (respresenting a default image) with $OR=0.21$ and $p<0.05$.
@@ 6.1. Default Images in Text-to-image Generation
6.2. Mechanisms and Characteristics of Default Images in Generative Models
@@
    Emergence of default images due to unknown inputs: TTI models default to certain visual mo (“dtifs and styles (“default images”) when they encounter inputs outside their training data.
@@ 6.3. Explainability and Alignment of Text-to-Image Generation
What does ‘absolute ownership’, ‘nallopand’, or a ‘protruding goofus’ look like? Humans could, after some thinking, come up with some image for these abstract concepts. However, machines struggle to produce a result (cf. Figure 1). This is not surprising, given that these terms were likely not part of the TTI models’ training data. While polysemia and ambiguity will likely continue to be an issue in the alignment of TTI models with human intent (consider, for instance, the prompt “a dog sitting in front of a bank”), default images are a different phenomenon – a type of image that appears in response to prompts that are not understood by the model.
@@ 7. Conclusion
We presented the first systematic investigation on default images in Midjourney. Default images are images with a certain motive and style that occur in response to a variety of different prompts. Through a structured prompt‐generation process and affinity diagramming, we identified ten canonical default images and distilled nine postulates characterizing their properties. Our user study demonstrated that default images may negatively influence user satisfaction. By defining the notion of default images, this work lays the groundwork for future work on the limits of text‐to‐image systems.

### v2

# An Exploration of Default Images in Text-to-Image Generation

@@ Abstract.
The figure shows different default images, resulting from unrelated input prompts. It is clearly visible that these images have a common theme, motif (i.e., subjects), and style.
@@ 1. Introduction
While there is a growing number of studies investigating TTI and the practice of prompt engineering, little attention has been paid to what happens when the model does not understand its inputs. TTI models are trained to always yield an output, even if the input prompt contains terms and concepts not known to the model. Therefore, when such a prompt is given to the TTI model, it may generate what we call default images – images that are seemingly unrelated to the input prompt and often do not match the user’s expectations. One key characteristic of default images is that the same (or very similar) default image may appear when prompted with different, unrelated prompts (see Figure 1 and Figure 3).
@@
In this paper, we present the first preliminary exploration into default images in Midjourney111www.midjourney.com, a popular image generation system. We describe our approach to systematically create a set of input prompts likely triggering default images, and describe our experiments in generating images from these prompts. We then apply affinity diagramming to identify similarities between the generated images, resulting in a set of ten default images. We describe these images and make nine postulates about their characteristics. We further conducted a user study investigating whether default images affect user satisfaction, and how they align with human perception and expectations. Note that due to the vastness of Midjourney’s latent space, this initial investigation is not exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and the importance of studying default images.
@@
  • A description of our systematic approach for generating default images. We contribute six categories of prompts that are likely to elicit default images as a valuable resource enabling future research on this topic.
@@
Due to the vastness of Midjourney’s latent space, this investigation can never be exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and present the results of several ablation studies.
2. Importance of Studying Default Images in Text-to-Image Generation
@@ 4.1. What are Default Images?
The defining characteristic of default images is that they are encountered in response to multiple different and—crucially—unrelated prompts. The examples given in Figure 1 and Figure 3 demonstrate how different unrelated prompts trigger very similar default images in Midjourney – images with a specific motif and style that originate from a region in the TTI model’s latent space that is not strongly associated with known concepts. Another defining characteristic of default images is that, while there are slight variations, the default images are remarkably stable in terms of their motif (i.e., subject) and style (cf. Figure 1). Their recurring motif and style makes default images distinctly recognizable.
@@ 5.1. Input Prompts Likely Triggering Default Images
In our initial exploration of the phenomenon of default images, we explored three categories of prompts that could result in a conceptual mismatch between the resulting image and user expectations. We first tried to ‘confuse’ the model with polysemous words or prompts outside the training data such as e-mail addresses, but we found the rate of default images too low for our purposes.
We then tried prompts that are likely not recognized by the model and lack practical purpose. These prompts include, but are not limited to: contextually misleading phrases (e.g., Einstein’s relativity theology), random character strings and special characters (e.g. aölkjnewjbtn and %), misheard lyrics, and abstract and metaphorical concepts that may not invoke strong imagery (such as absolute ownership and extreme aptitude). While prompts in this category may lead to default images, we excluded them from our analysis due to their limited relevance in real-world applications. These prompts do not reflect how users would typically interact with the TTI model, although they are relevant to for future investigations.
We then turned to create prompts that combine both not recognized terms and practical use cases, which we deem most promising to our endeavor. These prompts are likely used in the real world, but still challenge the model. This includes:
@@
The six categories of prompts emphasize real-world usefulness and they are our focus area and inspiration for creating default images. The low-resource languages were chosen for their language-native words that are not borrowed from other languages and, thus, do not cognate with widely known foreign languages. We focus on Finnish in this study (see Appendix A4.1), leaving Tagalog (Appendix A4.2) to future work. Using this approach, we created a set of prompts, listed in Appendix A. Each of the prompts was checked on haveibeentrained.com, a website that allows to search the training data used in many TTI models (including LAION (LAION.ai, nd; Schuhmann et al., 2022)). Prompts were only included if a search did not yield large numbers of images as a search result. We then generated images with these prompts on Midjourney, as described in the following section.
@@ 5.2. Image Generation Experiments and Ablation Studies
Our experiments were executed on Midjourney’s webpage222www.midjourney.com/imagine using fixed parameters listed in Table 1. For each of the six different categories of inputs A1–A6, we ran 20 trials (with exception of category glitch tokens, for which we ran 10 trials due to limited availability of glitch tokens), introducing slight variations between each run by corrupting or altering words (see the prompts in Appendix A). This resulted in 440 images. The images were then visually analyzed, as described in the following section.
@@ 5.4. Survey Study
In addition to our visual analysis with affinity diagramming, we utilized crowdsourcing on Prolific333www.prolific.com, an online human subject pool, to inquire about people’s perception of default images. In particular, default images may negatively affect user satisfaction when prompting TTI models. To investigate this, we designed an online study with a questionnaire in which four images were presented, including the prompt that created the image. Unknown to the participant, two of the images (Q1 and Q2) had one term in the prompt replaced with a term that caused default images in our experiments. For instance, in Q1, we replaced the term ‘raven’ in the prompt “photoshoot of a woman wearing a white off-shoulder dress with a black raven on top of her head” with the word ’greagoft’. The participant, however, was shown the original prompt and the image resulting from the modified prompt. This setup simulates the surprising mismatch in expectations that an encounter with a default image may produce in practice. For the third question (Q3), we used an image that matches exactly the descriptive prompt, without replacements. In the fourth question (Q4), we used a prompt which is unknown to the model and, thus, the accompanying image is a default image. The four prompts and images are depicted in Figure 6. Participants were asked to rate their satisfaction with each image, given the prompt, on a scale from 1 (Strongly Dissatisfied) to 7 (Strongly Satisfied). The reward to participants was £6.29 per hour, and the median time to complete the survey was 3 min and 49 seconds.
@@ 6.2.1. Effects of seed values
We visually analyzed how different seed values affect the default images. Using the different seeds with the same prompt resulted in different default images, partially sharing some of the motifs (see left images in Figure 5A). While the premise is still the same, one might not associate these images with each other. Thus, the best way to generate default images is by having a fixed seed value for all image generation runs. The images on the right of Figure 5A show how common elements may occur in default images. The commonalities in the default images included portraits of women and animals that were mostly birds.
Figure 5. Ablation studies: (A) Default images generated with different seed values (A1) and image similarities with variations in both seed values and prompts (A2). (B) The default images on the left demonstrate how known terms (computer, browsing) overpower the unknown terms (in bold), producing images that are not default images. The two images on the right demonstrate how motifs from default images can be created in specific contexts. (C) Default images generated using the prompt “hevonen” and seed 123456 with Midjourney model version 6.1, 6.0, 5.2, and 5.1. Similarities in default images point to similarities in model training data. The four different images generated in response to one prompt (on the left) are the result of Midjourney’s latent space sampling and internal image diversity mechanisms.
@@ 6.3.2. User study results
Figure 6 shows that unknown terms potentially triggering default images may negatively affect user satisfaction, but not in all cases. In Q1, the effect of replacing ‘raven’ with ‘greagoft’ resulted in only subtle changes to the resulting image. The image appeared almost as if a raven could be there. The participants’ satisfaction, therefore, was high ($Mean=5.2$, $SD=1.8$). In Q2, where ‘leaves and flowers’ was replaced with ‘greagoft’, the divergence between prompt and resulting image was more noticeable and satisfaction was lower ($Mean=2.7$, $SD=1.9$). A Fisher’s exact test indicates that the way participants perceived the change (subtle vs. noticeable) is significantly related to their satisfaction level ($OR=0.074$, $p<0.05$). The odds ratio of 0.074 indicates that dissatisfaction (1–3 on the Likert scale) is much less likely when the change is subtle compared to noticeable. In other words, participants were significantly more likely to be satisfied (5–7) when the change in the image was subtle (Q1) rather than noticeable (Q2).
For the third question (Q3), we did not switch or omit any words in the prompt, and the image was the exact result of the prompt. Participants were satisfied with the resulting image ($Mean=4.5$, $SD=1.9$). In the fourth question (Q4), we used a prompt known to trigger default images. Participants showed particularly strong dissatisfaction with this image ($Mean=2.6$, $SD=2.0$). Again, there was a significant difference in user satisfaction between Q3 (representing an exact match between prompt and image) and Q4 (respresenting a default image) with $OR=0.21$ and $p<0.05$.
@@ 7.1. Default Images in Text-to-image Generation
7.2. Characteristics of Default Images in Generative Models
@@
    Emergence of default images due to unknown inputs: TTI models default to certain visuals (“default images”) with distinct motif and styles when they encounter inputs outside their training data.
@@ 7.3. Explainability and Alignment of Text-to-Image Generation
What does ‘absolute ownership’, ‘nallopand’, or a ‘protruding goofus’ look like? Humans could, after some thinking, come up with some image for these abstract concepts. However, machines struggle to produce a result (cf. Figure 1). This is not surprising, given that these terms were likely not part of the TTI models’ training data. While polysemia and ambiguity will likely continue to be an issue in the alignment of TTI models with human intent (consider, for instance, the prompt “a crane in front of a bank”), default images are a different phenomenon – a type of image that appears in response to prompts that are not understood by the model.
@@ 8. Conclusion
We presented the first systematic investigation on default images in Midjourney. Default images are images with a certain motif and style that occur in response to a variety of different prompts. Through a structured prompt‐generation process and affinity diagramming, we identified ten canonical default images and distilled nine postulates characterizing their properties. Our user study demonstrated that default images may negatively influence user satisfaction. By defining the notion of default images, this work lays the groundwork for future work on the limits of text‐to‐image systems.